# AdvI2I: Adversarial Image Attack on Image-to-Image Diffusion models

### v2

@@ 5 Conclusion
In this work, we present AdvI2I, a novel framework designed to expose a vulnerability previously underexplored in I2I diffusion models. Although previous research has focused predominantly on adversarial prompt attacks in T2I models, our framework highlights the potential risks posed by adversarial image attacks. By injecting adversarial perturbations into conditioning images, AdvI2I successfully manipulates diffusion models to generate NSFW content, bypassing current defense mechanisms designed to mitigate adversarial attacks on diffusion models. Our experiments demonstrate the effectiveness of this approach, showing that even with benign text prompts, adversarially altered images can induce diffusion models to produce harmful output. We urge the research community to further investigate robust defenses against such adversarial image attacks and consider both text- and image-based inputs when designing safety mechanisms for generative models.
Appendix A Appendix
A.1 Evaluation of model transferability.
We evaluate the transferability of adversarial image attacks from the SDv1.5-Inpainting model to other versions of SD inpainting models (SDv2.0, SDv2.1, SDv3.0). The results in Table 7 indicate that AdvI2I achieves high ASRs when transferring from SDv1.5 to SDv2.0 and SDv2.1 (81.0% and 86.0%, respectively). Its performance drops significantly when transferred to SDv3.0, with an ASR of only 36.0%, since SDv3.0 applies Transformer rather than U-Net for diffusion noise prediction. This suggests that transferability is relatively strong across similar model architectures, such as different versions of SDv1.x and SDv2.x.
Additionally, no experiments were conducted to measure the transferability of the attacks to InstructionPix2Pix because its model architecture differs from that of the SD models. Furthermore, the training image resolution of InstructionPix2Pix is 256x256, whereas SD models struggle to achieve effective editing results at this resolution. Therefore, a direct transferability test between these models would not yield meaningful insights due to their structural and resolution differences.
@@ Appendix A Appendix
| SDv1.5-Inpainting | AdvI2I | 83.5% | 81.0% | 86.0% | 36.0% |
| Adaptive | 79.5% | 76.5% | 79.0% | 35.0% |
A.2 Ablation Studies

### v3

@@ 5 Conclusion
In this work, we introduce AdvI2I, a novel adversarial attack framework that reveals a previously underexplored vulnerability in I2I diffusion models. While prior research has primarily focused on adversarial prompt attacks, our study highlights the significant risks posed by adversarial image-based attacks. By injecting adversarial perturbations into conditioning images, AdvI2I effectively manipulates diffusion models to generate NSFW content, bypassing existing defense mechanisms designed to mitigate adversarial threats. Our experimental results demonstrate the effectiveness of this attack strategy, indicating that current defense mechanisms remain inadequate in addressing adversarial image attacks, underscoring the need for more robust safeguards. Given the increasing integration of I2I diffusion models in various applications, it is imperative for the research community to develop comprehensive security measures that address adversarial risks from both textual and image-based inputs. We urge further investigation into robust defense strategies, and ethical considerations in the deployment of diffusion models to mitigate potential misuse and enhance the safety of generative AI systems.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Configuration of the Safe Latent Diffusion (SLD)
We observe that even the ”Medium” strength setting of SLD can substantially degrade the quality of images generated during benign image editing tasks with I2I diffusion models. To address this issue and enhance compatibility with I2I diffusion models, we adjust the SLD configuration accordingly. Specifically, we set the guidance scale to 1000, the warmup step to 7, the threshold to 0.01, the momentum scale to 0.3, and to 0.4.
Appendix B Evaluation of Model Transferability
We evaluate the transferability of adversarial image attacks from the SDv1.5-Inpainting model to other versions of SD inpainting models (SDv2.0, SDv2.1, SDv3.0). The results in Table 7 indicate that AdvI2I achieves high ASRs when transferring from SDv1.5 to SDv2.0 and SDv2.1 (80.5% and 84.0%, respectively). Its performance drops significantly when transferred to SDv3.0, with an ASR of only 34.0%. We conjecture this is due to differences in training data: SDv3.0 is trained on the different dataset filtered to exclude explicit content, as noted in (Esser et al., 2024). This suggests that our attack can expose the risk when the I2I model has the inherent ability to generate NSFW images, but could fail otherwise. Therefore, a potential future direction to enhance model safety is to totally nullify the NSFW concept from the model by thoroughly cleaning the training data.
@@ Appendix B Evaluation of Model Transferability
| SDv1.5-Inpainting | AdvI2I | 82.5% | 80.5% | 84.0% | 34.0% |
| Adaptive | 78.5% | 73.5% | 77.5% | 33.0% |
We also evaluated AdvI2I-Adaptive under defenses across multiple I2I models. The results shown in Table 8 demonstrate the attack persistence when transferring from SDv1.5 to (black-box) SDv2.0 and SDv2.1.
| Source Model | Target Model | w/o Defense | SLD | SD-NP | GN | SC |
| SDv1.5-Inpainting | SDv1.5 | 78.5% | 75.0% | 75.5% | 72.5% | 72.0% |
| SDv2.0 | 73.5% | 72.5% | 75.5% | 69.5% | 67.0% | |
| SDv2.1 | 77.5% | 73.0% | 76.0% | 73.0% | 70.0% | |
| SDv3.0 | 33.0% | 30.5% | 30.5% | 27.0% | 30.0% |
Considering larger model difference, we evaluated the transferability from SDv1.5-Inpainting to FLUX.1-dev ControlNet Inpainting-Alpha and SDXL-Turbo. The results are shown in Table 9.
| Source Model | Target Model | ASR |
| SDv1.5-Inpainting | FLUX.1-dev ControlNet Inpainting-Alpha | 74.0% |
| SDv1.5-Inpainting | SDXL-Turbo | 62.5% |
Appendix C Ablation Studies
@@
Appendix D Results on the SDv2.1-Inpainting Model
We evaluate AdvI2I on the SDv2.1-Inpainting model. As shown in Table 12, it achieves an ASR of 78.5% under the nudity concept, demonstrating that AdvI2I can generalize to state-of-the-art diffusion models.
| Concept | Method | w/o Defense | SLD | SD-NP | GN | SC |
| Nudity | Attack VAE | 35.5% | 32.5% | 35.0% | 32.5% | 7.0% |
| MMA | 38.0% | 32.5% | 36.5% | 23.5% | 37.0% | |
| AdvI2I (ours) | 78.5% | 73.0% | 75.0% | 64.5% | 10.5% |
| Source Safety Checker | Target Safety Checke | ASR |
| ViT-L/14-based | ViT-L/14-based | 72.0% |
| ViT-B/32-based | 66.5% |
Appendix E The Transderability of AdvI2I-Adaptive on Differenet Safety Checkers
In our work, we consider a ViT-L/14-based NSFW-detector as the safety checker. We also evaluate the transferability of AdvI2I-Adaptive on SDv1.5-Inpainting to a ViT-B/32-based NSFW-detector and observe that it still achieves a high ASR, as shown in Table 13.
Appendix F The Evaluation of The Image Quality
We provide a comparison of the quality of attacked images using LPIPS, SSIM, PSNR, FSIM, and VIF. The results are in Table 14. The results highlight that AdvI2I performs on par with Attack VAE in terms of structural and perceptual similarity (SSIM and LPIPS) and visual feature retention (FSIM and VIF), while significantly outperforming MMA. Importantly, both AdvI2I and Attack VAE use generators to produce adversarial images, while MMA directly optimizes adversarial noise. Although MMA achieves a higher PSNR due to its direct noise optimization approach, it performs worse in metrics like VIF and SSIM. AdvI2I successfully balances adversarial effectiveness and attacked image quality across all metrics, reinforcing its stealthiness and robustness.
We include Face-Adapter (Han et al., 2025), a diffusion-based face swap method using SDv1.5 as the base model, as a baseline for comparison. The image quality is evaluated using multiple metrics: TOPIQ with three checkpoints trained on different datasets: flive, koniq, and spaq) (Chen et al., 2024), NIQE, PIQE, and FID. As shown in Table 15, AdvI2I consistently performs competitively across various metrics. It achieves higher quality in TOPIQ-koniq and TOPIQ-spaq compared to Face-Adapter, while also showing significant improvements in NIQE, PIQE, and FID scores, which indicate better perceptual quality and closer alignment to real image distributions. These results demonstrate that AdvI2I effectively generates high-quality adversarial images while maintaining its primary objective of exposing vulnerabilities in I2I models.
| Method | LPIPS↓ | SSIM↑ | PSNR↑ | FSIM↑ | VIF↑ | ASR(%)↑ |
| Attack VAE | 0.31 | 0.89 | 18.80 | 0.96 | 0.73 | 41.5 |
| MMA | 0.32 | 0.63 | 23.19 | 0.94 | 0.35 | 42.0 |
| AdvI2I (ours) | 0.31 | 0.88 | 18.79 | 0.96 | 0.72 | 82.5 |
| Method | TOPIQ-koniq↑ | TOPIQ-flive↑ | TOPIQ-spaq↑ | NIQE↓ | PIQE↓ | FID↓ |
| Face-Adapter | 0.43 | 0.83 | 0.50 | 6.36 | 62.60 | 104.63 |
| AdvI2I (ours) | 0.58 | 0.78 | 0.67 | 3.76 | 38.72 | 85.60 |
Appendix G Evaluation on more concepts
In addition to the ”nudity” and ”violence” concepts, we further evaluate the ”political extremism” concept. The concept vector is constructed with prompts related to ”extremism” and ”terrorism”. The results in Table 16 confirm AdvI2I’s versatility across diverse NSFW concepts.
| Method | Concept | w/o Defense | SLD | SD-NP | GN | SC |
| AdvI2I | Extremism | 76.5% | 73.0% | 73.5% | 60.5% | 27.5% |
| AdvI2I-Adaptive | Extremism | 74.5% | 70.0% | 72.5% | 71.5% | 72.0% |
Appendix H Robustness of AdvI2I against DiffPure
We evaluate the robustness of AdvI2I against DiffPure (Nie et al., 2022), a diffusion-based image purification defense. As shown in Table 17 When applied to the SDv1.5-Inpainting model on the nudity concept, DiffPure reduces the ASR of AdvI2I from 82.5% to 72.5%. This relatively small decrease suggests that AdvI2I is resilient to such purification-based defenses. We attribute this robustness to the fact that adversarial images in AdvI2I are generated via a learned generator, rather than being perturbed through additive noise.
| Method | w/o Defense | DiffPure |
| Attack VAE | 41.5% | 33.5% |
| AdvI2I (ours) | 82.5% | 72.5% |
Appendix I Exploring AdvI2I as a Defensive Mechanism
While the primary focus of this work is on attacking diffusion models via adversarial images, we also conduct a preliminary study to explore the potential of AdvI2I as a defensive mechanism.
Specifically, we investigate whether embedding a benign concept into an image—such as wearing clothes—can reduce the effectiveness of adversarial or explicit prompts during image generation. To this end, we use AdvI2I to embed the ”wearing clothes” concept into clean images, then evaluate how this affects the generation outcome when attacked with explicit prompts (e.g., “Make the woman naked”) using the SDv1.5-Inpainting model.
As shown in Table 18, embedding this benign concept reduces the ASR from 96.5% to 24.5%, suggesting that AdvI2I can be adapted as a conceptual defense to counter harmful generations.
| Input Condition | ASR on Explicit Prompt |
| Original Image | 96.5% |
| + AdvI2I (Wearing Clothes) | 24.5% |
These findings highlight the conceptual versatility of AdvI2I and motivate future work in leveraging image-conditioned generation methods as proactive defenses in diffusion models.